# Inference for Regression with

### v2

@@ A.2 Proofs for Section 3
$\left\|\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}(\hat{\bm{\theta}}_{i}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}-\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T})-(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\left(\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}(\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}-\mathbf{p}_{i}\!\!\!\;\mathbf{p}_{i}^{T})\right)\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}\right\|\to_{p}0,$
by Assumption 2(i)-(iii),(v). Since $\mathbf{p}_{i}$ and $\hat{\mathbf{p}}_{i}$ both take values in the simplex, we have $\|\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}-\mathbf{p}_{i}\!\!\!\;\mathbf{p}_{i}^{T}\|\leq 2\|\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}\|$ and so
$\left\|\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}(\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}-\mathbf{p}_{i}\!\!\!\;\mathbf{p}_{i}^{T})\right\|\leq 2\left(\max_{1\leq i\leq n}\|\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}\|\right)\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}\to_{p}0,$
@@
$T_{5,n}=\frac{2}{n}\sum_{i=1}^{n}(\hat{\varepsilon}_{i}-\varepsilon_{i})\varepsilon_{i}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}+\frac{1}{n}\sum_{i=1}^{n}(\hat{\varepsilon}_{i}-\varepsilon_{i})^{2}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T},$
@@
$=\left(\max_{1\leq i\leq n}|\hat{\varepsilon}_{i}-\varepsilon_{i}|\right)\mathrm{tr}\left\{\frac{2}{n}\sum_{i=1}^{n}|\varepsilon_{i}|\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\right\}+\left(\max_{1\leq i\leq n}|\hat{\varepsilon}_{i}-\varepsilon_{i}|^{2}\right)\mathrm{tr}\left\{\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\right\}\to_{p}0,$
because $\frac{1}{n}\sum_{i=1}^{n}|\varepsilon_{i}|\hat{\bm{\xi}}_{i}\!\!\!\!\;\hat{\bm{\xi}}_{i}^{T}$ is bounded in probability by control of $T_{3,n}$ and $T_{4,n}$, which together imply $\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}$ is bounded in probability, and $\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}$ is bounded in probability by Lemma 5. ∎
@@
$\mathbb{E}\left[\left.\hat{\mathbf{p}}_{t,i}^{T}\mathbf{M}\hat{\mathbf{p}}_{t,i}\right|F_{i}\right]=\mathbf{p}_{t,i}^{T}\mathbf{M}\mathbf{p}_{t,i}+\frac{1}{C_{t,i}}\,\mathrm{tr}\left\{\mathbf{M}\left(\mathrm{diag}(\mathbf{p}_{t,i})-\mathbf{p}_{t,i}\mathbf{p}_{t,i}^{T}\right)\right\}.$
@@
$\mathbb{E}\left[\left.\hat{S}_{i}^{2}\right|F_{i}\right]$ $=\mathbb{E}\left[\left.\mathbb{E}\left[\left.\hat{\mathbf{p}}_{1,i}^{T}(\hat{\mathbf{p}}_{2,i}\hat{\mathbf{p}}_{2,i}^{T})\hat{\mathbf{p}}_{1,i}\right|\hat{\mathbf{p}}_{2,i},F_{i}\right]\right|F_{i}\right]$
$=\mathbb{E}\left[\left.\mathbf{p}_{1,i}^{T}(\hat{\mathbf{p}}_{2,i}\hat{\mathbf{p}}_{2,i}^{T})\mathbf{p}_{1,i}+\frac{1}{C_{1,i}}\,\mathrm{tr}\left\{(\hat{\mathbf{p}}_{2,i}\hat{\mathbf{p}}_{2,i}^{T})\left(\mathrm{diag}(\mathbf{p}_{1,i})-\mathbf{p}_{1,i}\mathbf{p}_{1,i}^{T}\right)\right\}\right|F_{i}\right]$
$=\mathbb{E}\left[\left.\hat{\mathbf{p}}_{2,i}^{T}(\mathbf{p}_{1,i}\mathbf{p}_{1,i}^{T})\hat{\mathbf{p}}_{2,i}\right|F_{i}\right]+\frac{1}{C_{1,i}}\mathbb{E}\left[\left.\hat{\mathbf{p}}_{2,i}^{T}\left(\mathrm{diag}(\mathbf{p}_{1,i})-\mathbf{p}_{1,i}\mathbf{p}_{1,i}^{T}\right)\hat{\mathbf{p}}_{2,i}\right|F_{i}\right]$
$=S_{i}^{2}+\frac{1}{C_{2,i}}\mathbf{p}_{1,i}^{T}\left(\mathrm{diag}(\mathbf{p}_{2,i})-\mathbf{p}_{2,i}\mathbf{p}_{2,i}^{T}\right)\mathbf{p}_{1,i}+\frac{1}{C_{1,i}}\mathbf{p}_{2,i}^{T}\left(\mathrm{diag}(\mathbf{p}_{1,i})-\mathbf{p}_{1,i}\mathbf{p}_{1,i}^{T}\right)\mathbf{p}_{2,i}$
$\quad\quad+\frac{1}{C_{1,i}C_{2,i}}\mathrm{tr}\left\{\left(\mathrm{diag}(\mathbf{p}_{2,i})-\mathbf{p}_{2,i}\mathbf{p}_{2,i}^{T}\right)\left(\mathrm{diag}(\mathbf{p}_{1,i})-\mathbf{p}_{1,i}\mathbf{p}_{1,i}^{T}\right)\right\}.$ (27)
@@
$\mathbb{E}[\sqrt{n}(\hat{S}_{i}-S_{i})\hat{S}_{i}]\to\kappa_{1}\mathbb{E}\left[\mathbf{p}_{2,i}^{T}\left(\mathrm{diag}(\mathbf{p}_{1,i})-\mathbf{p}_{1,i}\mathbf{p}_{1,i}^{T}\right)\mathbf{p}_{2,i}\right]\\ +\kappa_{2}\mathbb{E}\left[\mathbf{p}_{1,i}^{T}\left(\mathrm{diag}(\mathbf{p}_{2,i})-\mathbf{p}_{2,i}\mathbf{p}_{2,i}^{T}\right)\mathbf{p}_{1,i}\right],$
@@ Appendix C Example Code
[…]
import numpyro.distributions as dist
import jax.numpy as jnp
from jax.nn import softmax
[…]
def __init__(self, K, N, V, z, q, eta = .1, alpha = 1):
self.K = K # number of latent types
self.N = N # number of observations
self.V = V # number of distinct features
self.z = z # number of covariates affecting outcome
self.q = q # number of covariates affecting type shares
self.eta = eta
self.alpha = alpha
[…]
# Supervised topic model with covariates
[…]
# X : feature count matrix
# C : total number of features per observation
# Z : covariates entering regression
# Q : covariates entering type shares
# K : number of types
# eta, alpha : Dirichlet hyperparamters
[…]
beta = sample(”beta”, dist.Dirichlet(
self.eta * jnp.ones(self.V - self.num_anchors_per_class)))
[…]
A = sample(”A”, dist.Normal(jnp.matmul(Q, phis) , self.alpha))
[…]
theta = deterministic(
%**** 0file.tex Line 1775 ****”theta”,
softmax(jnp.hstack([A, jnp.zeros([self.D, 1])]), axis = -1)
[…]
total_count=C,
probs = jnp.matmul(theta, beta)
with plate(”hist”, self.N):
X_bows = sample(”obs_x”, distMultinomial, obs = X)
[…]
zetas = sample(”zetas”, dist.Normal(0, 10).expand([self.z]))
sigma = sample(”sigma”, dist.Gamma(1, 10))
[…]
Y = sample(”obs_y”, dist.Normal(mean, sigma), obs = Y)

### v3

@@ A.2 Proofs for Section 3
$\left\|\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}(\hat{\bm{\theta}}_{i}^{\phantom{T}}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}-\bm{\theta}_{i}^{\phantom{T}}\!\!\!\;\bm{\theta}_{i}^{T})-(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\left(\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}(\hat{\mathbf{p}}_{i}^{\phantom{T}}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}-\mathbf{p}_{i}^{\phantom{T}}\!\!\!\;\mathbf{p}_{i}^{T})\right)\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}\right\|\to_{p}0,$
by Assumption 2(i)-(iii),(v). Since $\mathbf{p}_{i}$ and $\hat{\mathbf{p}}_{i}$ both take values in the simplex, we have $\|\hat{\mathbf{p}}_{i}^{\phantom{T}}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}-\mathbf{p}_{i}^{\phantom{T}}\!\!\!\;\mathbf{p}_{i}^{T}\|\leq 2\|\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}\|$ and so
$\left\|\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}(\hat{\mathbf{p}}_{i}^{\phantom{T}}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}-\mathbf{p}_{i}^{\phantom{T}}\!\!\!\;\mathbf{p}_{i}^{T})\right\|\leq 2\left(\max_{1\leq i\leq n}\|\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}\|\right)\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}\to_{p}0,$
@@
$T_{5,n}=\frac{2}{n}\sum_{i=1}^{n}(\hat{\varepsilon}_{i}-\varepsilon_{i})\varepsilon_{i}\hat{\bm{\xi}}_{i}^{\phantom{T}}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}+\frac{1}{n}\sum_{i=1}^{n}(\hat{\varepsilon}_{i}-\varepsilon_{i})^{2}\hat{\bm{\xi}}_{i}^{\phantom{T}}\!\!\!\;\hat{\bm{\xi}}_{i}^{T},$
@@
$=\left(\max_{1\leq i\leq n}|\hat{\varepsilon}_{i}-\varepsilon_{i}|\right)\mathrm{tr}\left\{\frac{2}{n}\sum_{i=1}^{n}|\varepsilon_{i}|\hat{\bm{\xi}}_{i}^{\phantom{T}}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\right\}+\left(\max_{1\leq i\leq n}|\hat{\varepsilon}_{i}-\varepsilon_{i}|^{2}\right)\mathrm{tr}\left\{\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}^{\phantom{T}}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\right\}\to_{p}0,$
because $\frac{1}{n}\sum_{i=1}^{n}|\varepsilon_{i}|\hat{\bm{\xi}}_{i}^{\phantom{T}}\!\!\!\!\;\hat{\bm{\xi}}_{i}^{T}$ is bounded in probability by control of $T_{3,n}$ and $T_{4,n}$, which together imply $\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}\hat{\bm{\xi}}_{i}^{\phantom{T}}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}$ is bounded in probability, and $\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}^{\phantom{T}}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}$ is bounded in probability by Lemma 5. ∎
@@
$\mathbb{E}\left[\left.\hat{\mathbf{p}}_{t,i}^{T}\mathbf{M}\hat{\mathbf{p}}_{t,i}^{\phantom{T}}\right|F_{i}\right]=\mathbf{p}_{t,i}^{T}\mathbf{M}\mathbf{p}_{t,i}^{\phantom{T}}+\frac{1}{C_{t,i}}\,\mathrm{tr}\left\{\mathbf{M}\left(\mathrm{diag}(\mathbf{p}_{t,i})-\mathbf{p}_{t,i}^{\phantom{T}}\mathbf{p}_{t,i}^{T}\right)\right\}.$
@@
$\mathbb{E}\left[\left.\hat{S}_{i}^{2}\right|F_{i}\right]$ $=\mathbb{E}\left[\left.\mathbb{E}\left[\left.\hat{\mathbf{p}}_{1,i}^{T}(\hat{\mathbf{p}}_{2,i}^{\phantom{T}}\hat{\mathbf{p}}_{2,i}^{T})\hat{\mathbf{p}}_{1,i}^{\phantom{T}}\right|\hat{\mathbf{p}}_{2,i},F_{i}\right]\right|F_{i}\right]$
$=\mathbb{E}\left[\left.\mathbf{p}_{1,i}^{T}(\hat{\mathbf{p}}_{2,i}^{\phantom{T}}\hat{\mathbf{p}}_{2,i}^{T})\mathbf{p}_{1,i}^{\phantom{T}}+\frac{1}{C_{1,i}}\,\mathrm{tr}\left\{(\hat{\mathbf{p}}_{2,i}^{\phantom{T}}\hat{\mathbf{p}}_{2,i}^{T})\left(\mathrm{diag}(\mathbf{p}_{1,i})-\mathbf{p}_{1,i}^{\phantom{T}}\mathbf{p}_{1,i}^{T}\right)\right\}\right|F_{i}\right]$
$=\mathbb{E}\left[\left.\hat{\mathbf{p}}_{2,i}^{T}(\mathbf{p}_{1,i}^{\phantom{T}}\mathbf{p}_{1,i}^{T})\hat{\mathbf{p}}_{2,i}^{\phantom{T}}\right|F_{i}\right]+\frac{1}{C_{1,i}}\mathbb{E}\left[\left.\hat{\mathbf{p}}_{2,i}^{T}\left(\mathrm{diag}(\mathbf{p}_{1,i})-\mathbf{p}_{1,i}^{\phantom{T}}\mathbf{p}_{1,i}^{T}\right)\hat{\mathbf{p}}_{2,i}^{\phantom{T}}\right|F_{i}\right]$
$=S_{i}^{2}+\frac{1}{C_{2,i}}\mathbf{p}_{1,i}^{T}\left(\mathrm{diag}(\mathbf{p}_{2,i})-\mathbf{p}_{2,i}^{\phantom{T}}\mathbf{p}_{2,i}^{T}\right)\mathbf{p}_{1,i}^{\phantom{T}}+\frac{1}{C_{1,i}}\mathbf{p}_{2,i}^{T}\left(\mathrm{diag}(\mathbf{p}_{1,i})-\mathbf{p}_{1,i}^{\phantom{T}}\mathbf{p}_{1,i}^{T}\right)\mathbf{p}_{2,i}^{\phantom{T}}$
$\quad\quad+\frac{1}{C_{1,i}C_{2,i}}\mathrm{tr}\left\{\left(\mathrm{diag}(\mathbf{p}_{2,i})-\mathbf{p}_{2,i}^{\phantom{T}}\mathbf{p}_{2,i}^{T}\right)\left(\mathrm{diag}(\mathbf{p}_{1,i})-\mathbf{p}_{1,i}^{\phantom{T}}\mathbf{p}_{1,i}^{T}\right)\right\}.$ (28)
@@
$\mathbb{E}[\sqrt{n}(\hat{S}_{i}-S_{i})\hat{S}_{i}]\to\kappa_{1}\mathbb{E}\left[\mathbf{p}_{2,i}^{T}\left(\mathrm{diag}(\mathbf{p}_{1,i})-\mathbf{p}_{1,i}^{\phantom{T}}\mathbf{p}_{1,i}^{T}\right)\mathbf{p}_{2,i}^{\phantom{T}}\right]\\ +\kappa_{2}\mathbb{E}\left[\mathbf{p}_{1,i}^{T}\left(\mathrm{diag}(\mathbf{p}_{2,i})-\mathbf{p}_{2,i}^{\phantom{T}}\mathbf{p}_{2,i}^{T}\right)\mathbf{p}_{1,i}^{\phantom{T}}\right],$
@@ Appendix C Example Code
{minted}
[frame=lines,framesep=2mm,baselinestretch=.8,bgcolor=mygray,fontsize=,linenos]python
from numpyro import sample, plate import numpyro.distributions as dist import jax.numpy as jnp from jax.nn import softmax
class SUPTMC: def __init__(self, K, N, V, z, q, eta = .1, alpha = 1): self.K = K # number of latent types self.N = N # number of observations self.V = V # number of distinct features self.z = z # number of covariates affecting outcome self.q = q # number of covariates affecting type shares self.eta = eta self.alpha = alpha
def model(self, C, Z, Q, Y=None, X=None): # Supervised topic model with covariates
# Y : regression outcome # X : feature count matrix # C : total number of features per observation # Z : covariates entering regression # Q : covariates entering type shares # K : number of types # eta, alpha : Dirichlet hyperparamters
#### Upstream Factor Model ###
with plate(”topics”, self.K): beta = sample(”beta”, dist.Dirichlet( self.eta * jnp.ones(self.V - self.num_anchors_per_class)))
phis = sample(”phis”, dist.Normal(0,2).expand([self.q, self.K-1]))
with plate_stack(”docs”, sizes = [self.N, self.K - 1]): A = sample(”A”, dist.Normal(jnp.matmul(Q, phis) , self.alpha))
# document-topic distributions theta = deterministic( ”theta”, softmax(jnp.hstack([A, jnp.zeros([self.D, 1])]), axis = -1) )
distMultinomial = dist.Multinomial( total_count=C, probs = jnp.matmul(theta, beta) ) with plate(”hist”, self.N): X_bows = sample(”obs_x”, distMultinomial, obs = X)
#### Downstream Regression Model ###
gammas = sample(”gammas”, dist.Normal(0, 10).expand([self.K-1])) zetas = sample(”zetas”, dist.Normal(0, 10).expand([self.z])) sigma = sample(”sigma”, dist.Gamma(1, 10))
mean = jnp.matmul(theta[:,:(self.K-1)], gammas) + jnp.matmul(Z, zetas)
with plate(”y”, self.N): Y = sample(”obs_y”, dist.Normal(mean, sigma), obs = Y)